OpenStudy (anonymous):

A group of students equally shouldered the price of a broken laboratory instrument which is 60 dollars.If there had been three more students in their group, the cost of each share would have been a dollar less.How many students were in the group? (i need the equation also)

OpenStudy (anonymous):

Let there be x students in the group. Then their share is 60/x. If there were (x+3) students their share would be 60/(x+3). According to the problem, 60/x = 60/(x+3) + 1 or, 60/x = (x + 63)/(x + 3) or, 60x + 180 = x² + 63x or, x² + 3x - 180 = 0 It solves to x = 12. So, there were 12 students.